OpenStudy (anonymous):

Help will give medal and fan whoevr helps...Brooklyn opened a store credit card to purchase a sofa for $849. She put the entire purchase on the credit card. Her APR is 19.99% and the minimum payment is 5%. Assuming she makes only minimum payments, what is the remaining balance on her card after 3 months? $727.91 $741.31 $820.69 $766.88

OpenStudy (anonymous):

1st month 849 x (0.1999 / 12) = $14.14 interest 849 + 14.14 = 863.14 balance before payment 863.14 x 0.05 = 43.16 Payment 863.14 - 43.16 = 819.98 balance after payment 2nd month 819.98 x (0.1999 / 12) = $13.66 interest 819.98 + 13.66 = 833.64 balance before payment 833.64 x 0.05 = 41.68 Payment 833.64 - 41.68 = 791.96 balance after payment 3rd month 791.96 x (0.1999 / 12) = $13.19 interest 791.96 + 13.19 = 805.15 balance before payment 805.15 x 0.05 = 40.26 Payment 805.15 - 40.26 = 764.89 balance after payment So the answer is $766.88 which is closest to what we have.

OpenStudy (anonymous):

1st month 849 x (0.1999 / 12) = $14.14 interest 849 + 14.14 = 863.14 balance before payment 863.14 x 0.05 = 43.16 Payment 863.14 - 43.16 = 819.98 balance after payment 2nd month 819.98 x (0.1999 / 12) = $13.66 interest 819.98 + 13.66 = 833.64 balance before payment 833.64 x 0.05 = 41.68 Payment 833.64 - 41.68 = 791.96 balance after payment 3rd month 791.96 x (0.1999 / 12) = $13.19 interest 791.96 + 13.19 = 805.15 balance before payment 805.15 x 0.05 = 40.26 Payment 805.15 - 40.26 = 764.89 balance after payment

OpenStudy (anonymous):

will somone support @Fall12-13 anser

OpenStudy (anonymous):

thanks @CourageTheCowardlyDog but it doesnt give me any of the multiple choices above

OpenStudy (anonymous):

anyone?

jimthompson5910 (jim_thompson5910):

I'm using this calculator http://www.calculator.net/credit-card-calculator.html?balance=849&rate=19.99&minimum=42.45&payoffoption=0&fixedpaymentamount=35&x=67&y=22# but I'm getting $762.66 So it's a bit off

OpenStudy (anonymous):

@jim_thompson5910 So either way, aren't we supposed to take the closest value?

jimthompson5910 (jim_thompson5910):

yeah but I was hoping the formula/calculator would lead to one of the exact choices